OpenStudy (anonymous):

(Linear Equations) Find the slope of the line through the points (-3, 2) and (2, 3).

OpenStudy (watchmath):

Just plugin to the formula \[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

5/5=5

OpenStudy (anonymous):

5/1=5

OpenStudy (watchmath):

Be careful .. \(m=\frac{3-2}{2-(-3)}=\frac{1}{5}\).